OpenStudy (anonymous):

Help Please ?! [02.05]Solve for x: 3 − (2x − 5) < −4(x + 2)

OpenStudy (anonymous):

Anyone please :/

TheSmartOne (thesmartone):

Hint: \(\sf\Large a(b+c) = ab+ac\) \(\sf\Large a(b-c)=ab-ac\)

TheSmartOne (thesmartone):

also, hint: \(\sf\Large -a(b+c)=-ab-ac\) \(\sf\Large -a(b-c) = -ab+ac\)

OpenStudy (anonymous):

x<−8